# Carrier Aggregation Enabled MIMO-OFDM Integrated Sensing and Communication

### v1

@@ Abstract
In the evolution towards the forthcoming era of sixth-generation (6G) mobile communication systems characterized by ubiquitous intelligence, integrated sensing and communication (ISAC) is in a phase of burgeoning development. However, the capabilities of communication and sensing within single frequency band fall short of meeting the escalating demands. To this end, this paper introduces a carrier aggregation (CA)-enabled multi-input multi-output orthogonal frequency division multiplexing (MIMO-OFDM) ISAC system fusing the sensing data on high and low-frequency bands by symbol-level fusion for ultimate communication experience and high-accuracy sensing. The challenges in sensing signal processing introduced by CA include the initial phase misalignment of the echo signals on high and low-frequency bands due to attenuation and radar cross section, and the fusion of the sensing data on high and low-frequency bands with different physical-layer parameters. To this end, the sensing signal processing is decomposed into two stages. In the first stage, the problem of initial phase misalignment of the echo signals on high and low-frequency bands is solved by the angle compensation, space-domain diversity and vector cross-correlation operations. In the second stage, this paper realizes symbol-level fusion of the sensing data on high and low-frequency bands through sensing vector rearrangement and cyclic prefix adjustment operations, thereby obtaining high-precision sensing performance. Then, the closed-form communication mutual information (MI) and sensing Cramér-Rao lower bound (CRLB) for the proposed ISAC system are derived to explore the theoretical performance bound with CA. Simulation results validate the feasibility and superiority of the proposed ISAC system.
@@
Carrier aggregation (CA), integrated sensing and communication (ISAC), multi-band cooperative ISAC, multi-input multi-output (MIMO), orthogonal frequency division multiplexing (OFDM).
@@ I Introduction
The next-generation mobile communication system is anticipated to offer a plethora of multi-dimensional services extending beyond mere communication, such as smart transportation, smart factories, and digital twins [1], which have raised the demands for high-speed communication and high-accuracy sensing. To this end, integrated sensing and communication (ISAC) is considered to provide the high-speed communication experience and high-accuracy sensing [2, 3, 4, 5].
Nevertheless, the limitations imposed by sensing under single frequency band have spurred the study of ISAC over multiple frequency bands with the technique of carrier aggregation (CA). CA-enabled ISAC systems have the following potential advantages.
  • Communication: CA enhances bandwidth utilization and boosts communication transmission data rate while supporting various wireless access technologies, thereby improving overall communication performance [6].
@@
Introducing the concept of steeped-carrier OFDM
@@
Proposing a novel DFT method to address the
@@ A BWE method is proposed to recovery
Extending the BWE method to polarimetric radar
@@
  • Intra-band CA: Two primary ISAC signals are designed for intra-band CA: steeped-carrier OFDM ISAC signal and non-continuous (NC) fragmented OFDM ISAC signals. In terms of steeped-carrier OFDM ISAC signals, Pfeffer et al. [8] proposed the concept of steeped-carrier OFDM signals, which assigns different carrier frequencies to each resource block group during the continuous OFDM symbol times, expanding the overall bandwidth and enhancing the sensing performance. However, it fails to consider the Doppler frequency shift induced by rapidly moving targets. To this end, Schweizer et al. [9] explored a modified DFT method to mitigate the phase error caused by the high mobility of targets in steeped-carrier OFDM scheme, which obtains a high-accuracy velocity estimation. In terms of NC fragmented OFDM ISAC signals, Huang et al. [10] introduced a NC fragmented OFDM ISAC system, which can flexibly allocate spectrum bands and adapt to dynamic spectrum environment. However, the missing spectrum bands deteriorate Fourier sidelobes in target sensing. To address this problem, Liu et al. [11] proposed a joint compressed sensing (CS) and machine learning sensing method for fragmented spectrum bands to achieve high-accuracy estimation of range and velocity with low sidelobes.
@@
  • Inter-band CA: Multi-band radar stands as a representative example in inter-band CA. Cuomo [12] discussed a NC multi-band coherent radar and proposed a bandwidth extrapolation (BWE) method to recover missing spectrum bands, thereby enhancing the range resolution of radar. Suwa et al. [13] extended the BWE method to polarimetric radar, which achieves high resolution of range estimation.
@@
In summary, intra-band and inter-band CA enhance the efficient utilization of fragmented spectrum bands, leading to performance gains in both communication and sensing. Currently, the fragmented spectrum bands in mobile communication system are mostly dispersed, indicating broader applications in mobile communication system under inter-band CA. However, the investigation into inter-band CA in ISAC mobile communication systems is notably scarce. In our previous work, we have investigated the CA-enabled OFDM ISAC system, corroborating the effectiveness of CA through the validation [7]. However, our initial exploration, although valuable, remains somewhat rudimentary and idealized, overlooking the typical MIMO-OFDM signal in mobile communication systems.
@@
Hence, this paper investigates a CA-enabled MIMO-OFDM ISAC system where the fragmented high and low-frequency bands are shared for downlink (DL) communication and potential target sensing. In addition, a CA-enabled MIMO-OFDM ISAC signal model and the corresponding sensing processing method are presented. Through theoretical derivations and numerical simulations, we validate the superiority of the proposed ISAC system over conventional ISAC systems. The main contributions of this paper are summarized as follows.
@@
  • In terms of sensing processing, we have delved into the data-level and symbol-level fusion of the sensing data on high and low-frequency bands. In the symbol-level fusion of sensing data, two challenges manifest: 1) The initial phase misalignment of the echo signals on high and low-frequency bands due to attenuation and radar cross section (RCS); 2) The fusion of the sensing data on high and low-frequency bands with different physical-layer parameters. To this end, this paper decomposes the sensing signal processing into signal preprocessing stage and sensing information fusion stage. The signal preprocessing stage fully leverage the sensing data from multiple antennas while aligning the initial phases of the echo signals on high and low-frequency bands using the cyclic cross-correlation (CCC) method, as detailed in Section III-A. In the sensing information fusion stage, this paper realizes the symbol-level fusion of the sensing data on high and low-frequency bands, as well as the range and velocity estimations of target by rearranging the high and low-frequency feature vectors and using the premise of cyclic prefix (CP) dynamic adjustment, as detailed in Section III-B.
  • The theoretical boundaries of the proposed CA-enabled MIMO-OFDM ISAC signal are derived. Specifically, mutual information (MI) is adopted as the performance metric for communication, while the Cramér-Rao lower bound (CRLB) is utilized as a typical performance metric of sensing. Based on the proposed ISAC system, we derived the closed-form communication MI under the frequency-selective fading MIMO-OFDM channel model and sensing CRLBs for range and velocity estimations under the MIMO-OFDM sensing signal model.
@@ II System Model
We consider a CA-enabled MIMO-OFDM ISAC base station (BS) for DL communication and target sensing as shown in Fig. 1. The ISAC BS is equipped with $N_{\text{T}}+N_{\text{R}}$ antennas, where $N_{\text{T}}$ transmit antennas in transmitter (Tx) simultaneously serve $U$ user equipments (UEs) with $N_{\text{U}}$ antennas and detect potential targets. $N_{\text{R}}$ receive antennas in receiver (Rx) are equipped to receive the echo signals reflected by targets [14]. The spacing of antennas in BS is denoted by $d_{\text{r}}$ [15].
@@ II-A CA-enabled MIMO-OFDM ISAC Signal
represents the data for all UEs on $N_{b}$ subcarriers during $M_{b}$ OFDM symbol times, where $\mathbf{S}_{n}^{b}=\left[\mathbf{s}_{n,1}^{b},\mathbf{s}_{n,2}^{b},\cdots,\mathbf{s}_{n,U}^{b}\right]\in\mathbb{C}^{M_{b}\times U}$ indicates the data for all UEs on the $n$-th subcarrier during $M_{b}$ symbol times with $\mathbf{s}_{n,u}^{b}\in\mathbb{C}^{M_{b}\times 1}$ being the specific symbol data vector for the $u\in\{1,2,\cdots,U\}$-th UE [16]. To guarantee that the communication beam focuses on UEs and potential targets, the symbol data matrix undergoes a precoding operation [17]. Specifically, a compact precoding matrix [16]
@@
$x^{b}\left(k,n^{\prime},m\right)=\frac{1}{N_{b}}\sum_{n=0}^{N_{b}-1}\mathbf{X}_{n}^{b}\left(k,m\right)e^{j2\pi\frac{n^{\prime}n}{N_{b}}},$ (4)
where $\mathbf{X}_{n}^{b}\left(k,m\right)$ denotes the $\left(k,m\right)$-th element of $\mathbf{X}_{n}^{b}$, $n^{\prime}\in\{0,1,\cdots,N_{b}-1\}$ represents the index of subcarriers in time-domain and $k\in\{0,1,\cdots,N_{\text{T}}-1\}$ represents the index of transmit antenna. A CP with size $N_{\text{s}}^{b}$ greater than the maximum delay caused by target reflection is added to avoid inter-symbol interference. After digital-to-analog conversion (DAC), the analog signals of the $1$-st and $2$-nd CCs are up-converted using different local oscillators (LOs), mixed, and transmitted via a high power amplifier (HPA).
@@
$x(t)_{k}=\sum_{b=1}^{\mathcal{B}}\sum_{m=0}^{M_{b}-1}\sum_{n=0}^{N_{b}-1}\mathbf{X}_{n}^{b}\left(k,m\right)e^{j2\pi t(f_{\mathrm{C}}^{b}+n\Delta f^{b})}\text{rect}\left(\frac{t-mT^{b}}{T^{b}}\right),$ (5)
@@ II-B Communication Model
In the communication channel model, the differences in attenuations caused by high and low-frequency bands are not considered.
@@ II-B1 Communication channel model
is the channel matrix of the $l\in\{0,1,\cdots,L-1\}$-th delay path for the $u$-th UE with $h_{k,z}^{u,l,b}\sim\mathcal{CN}(0,1)$ being the corresponding channel element with independent identically distributed (i.i.d.) Rayleigh fading coefficients [16]. Therefore, the MIMO communication channel matrix in time-domain for the $u$-th UE on the $n$-th subcarrier during the $m$-th OFDM symbol time is expressed as [18]
@@ II-B2 Received communication signal
At the DL preamble information period, pilot preambles are used to obtain channel state information (CSI) estimation [19]. At the DL data information period, the $u$-th UE receives the DL communication signal, and undergoes down-conversion, analog-to-digital conversion (ADC), CP removal, discrete Fourier transform (DFT), and other operations to get a received data $\mathbf{Y}_{u}^{\mathrm{C}}$, which is expressed as
@@
In terms of communication performance in CA-enabled MIMO-OFDM ISAC system, the closed-form communication MI is derived in Section IV-A.
@@ II-C Sensing Model
In the sensing channel model, the differences in attenuations caused by high and low-frequency bands are considered. Then, the sensing channel model of the $b$-th CC and received sensing signal are derived.
@@ II-C1 Sensing channel model
We assume that the radial velocity of a potential target is $v_{0}$ and the range to BS is $r_{0}$. Meanwhile, the angle of arrive (AoA) and the angle of departure (AoD) between the target and BS are identical and are denoted by $\theta_{\text{Rx}}=\theta_{\text{Tx}}$. The transmit and received steering vectors are expressed as (9) and (10), respectively.
$\mathbf{a}_{\text{Rx}}(\theta_{\text{Rx}})$ $=\left[e^{j2\pi p(\frac{d_{\text{r}}}{\lambda^{b}})\sin(\theta_{\text{Rx}})}\right]^{\text{T}}|_{p=0,1,\cdots,N_{\text{R}}-1},$ (9)
$\mathbf{a}_{\text{Tx}}(\theta_{\text{Tx}})$ $=\left[e^{j2\pi k(\frac{d_{\text{r}}}{\lambda^{b}})\sin(\theta_{\text{Tx}})}\right]^{\text{T}}|_{k=0,1,\cdots,N_{\text{T}}-1},$ (10)
@@
The sensing channel model on the $n$-th subcarrier during the $m$-th OFDM symbol time in the $b$-th CC is expressed as
$\mathbf{H}_{m,n}^{b}=\kappa_{S}^{b}e^{j2\pi\left(f_{\text{s}}^{b}mT^{b}-n\Delta f^{b}\tau_{0}\right)}\mathbf{a}_{\text{Rx}}(\theta_{\text{Rx}})\mathbf{a}_{\text{Tx}}^{\text{T}}(\theta_{\text{Tx}}),$ (11)
where $\mathbf{H}_{m,n}^{b}\in\mathbb{C}^{N_{\text{R}}\times N_{\text{T}}}$ and $\kappa_{S}^{b}=\sqrt{\frac{(\lambda^{b})^{2}}{(4\pi)^{3}r_{0}^{4}}}\beta_{S}^{b}$ is the attenuation between the target and BS with $\beta_{S}^{b}\sim\mathcal{CN}(0,\sigma_{\beta}^{2})$ being the target RCS [20]; $f_{\text{s}}^{b}=\frac{2f_{\mathrm{C}}^{b}v_{0}}{c}$ and $\tau_{0}=\frac{2r_{0}}{c}$ are the Doppler frequency shift and delay, respectively.

### v2

@@ Abstract
In the evolution towards the forthcoming era of sixth-generation (6G) mobile communication systems characterized by ubiquitous intelligence, integrated sensing and communication (ISAC) is in a phase of burgeoning development. However, the capabilities of communication and sensing within single frequency band fall short of meeting the escalating demands. To this end, this paper introduces a carrier aggregation (CA)-enabled multi-input multi-output orthogonal frequency division multiplexing (MIMO-OFDM) ISAC system fusing the sensing data on high and low-frequency bands by symbol-level fusion for ultimate communication experience and high-accuracy sensing. The challenges in sensing signal processing introduced by CA include the initial phase misalignment of the echo signals on high and low-frequency bands due to attenuation and radar cross section, and the fusion of the sensing data on high and low-frequency bands with different physical-layer parameters. To this end, the sensing signal processing is decomposed into two stages. In the first stage, the problem of initial phase misalignment of the echo signals on high and low-frequency bands is solved by the angle compensation, spatial filtering and cyclic cross-correlation operations. In the second stage, this paper realizes symbol-level fusion of the sensing data on high and low-frequency bands through sensing vector rearrangement and cyclic prefix adjustment operations, thereby obtaining high-precision sensing performance. Then, the closed-form communication mutual information (MI) and sensing Cramér-Rao lower bound (CRLB) for the proposed ISAC system are derived to explore the theoretical performance bound with CA. Simulation results validate the feasibility and superiority of the proposed ISAC system.
@@
Carrier aggregation (CA), integrated sensing and communication (ISAC), multi-band cooperative ISAC, multi-input multi-output (MIMO), orthogonal frequency division multiplexing (OFDM), symbol-level fusion.
@@ I Introduction
The next-generation mobile communication system is anticipated to offer a plethora of multi-dimensional services extending beyond mere communication, such as smart transportation, smart factories, and digital twins [1, 2, 3, 4], which have raised the demands for high-speed communication and high-accuracy sensing. To this end, integrated sensing and communication (ISAC) is considered to provide the high-speed communication experience and high-accuracy sensing [5, 6, 7, 3].
Nevertheless, the limitations imposed by sensing under single frequency band have spurred the study of ISAC over multiple frequency bands with the technique of carrier aggregation (CA) [8, 3, 9, 10]. CA-enabled ISAC systems have the following potential advantages.
  • Communication: CA enhances bandwidth utilization and boosts communication transmission data rate while supporting various wireless access technologies, thereby improving overall communication performance [11, 8].
@@
Introduce the concept of stepped-carrier OFDM
@@
Propose a novel DFT method to address the
@@ A BWE method is proposed to recovery
Extend the BWE method to polarimetric radar
@@
  • Intra-band CA: Two primary ISAC signals are designed for intra-band CA: stepped-carrier OFDM ISAC signals and non-continuous (NC) fragmented OFDM ISAC signals. In terms of stepped-carrier OFDM ISAC signals, Pfeffer et al. in [12] proposed the concept of stepped-carrier OFDM signals, which assigns different carrier frequencies to each resource block group during the continuous OFDM symbol periods, expanding the overall bandwidth and enhancing the sensing performance. However, it fails to consider the Doppler frequency shift induced by rapidly moving targets. To this end, Schweizer et al. in [13] explored a modified DFT method to mitigate the phase error caused by the high mobility of targets in stepped-carrier OFDM scheme, which obtains a high-accuracy velocity estimation. In terms of NC fragmented OFDM ISAC signals, Huang et al. in [14] introduced a NC fragmented OFDM ISAC system, which can flexibly allocate spectrum bands and adapt to dynamic spectrum environment. However, the missing spectrum bands deteriorate Fourier sidelobes in target sensing. To address this problem, Liu et al. in [15] proposed a joint compressed sensing (CS) and machine learning sensing method for fragmented spectrum bands to achieve high-accuracy estimations of range and velocity with low sidelobes.
@@
  • Inter-band CA: Multi-band radar stands as a representative example in inter-band CA. Cuomo in [16] discussed a NC multi-band coherent radar and proposed a bandwidth extrapolation (BWE) method to recover missing spectrum bands, thereby enhancing the range resolution of radar. Suwa et al. in [17] extended the BWE method to polarimetric radar, which achieves high resolution of range estimation.
@@
In summary, intra-band and inter-band CA enhance the efficient utilization of fragmented frequency bands, leading to performance gains in both communication and sensing. Currently, the fragmented frequency bands in mobile communication system are mostly dispersed, indicating broader applications in mobile communication system under inter-band CA. However, the investigation into inter-band CA in ISAC mobile communication systems is notably scarce. In our previous work, we have investigated the CA-enabled OFDM ISAC system, corroborating the effectiveness of CA through the validation [8]. However, our initial exploration, although valuable, remains somewhat rudimentary and idealized, overlooking the typical MIMO-OFDM signal in mobile communication systems and the challenges of sensing signal processing caused by practical channel environment. Therefore, it is particularly crucial to further study the feasibility and practicability of inter-band CA-enabled ISAC signal processing.
@@
Hence, this paper investigates a CA-enabled MIMO-OFDM ISAC system where the fragmented high and low-frequency bands are shared for downlink (DL) communication and sensing potential targets. In addition, a CA-enabled MIMO-OFDM ISAC signal model and the corresponding sensing processing method are presented. Through theoretical derivations and numerical simulations, we validate the superiority of the proposed ISAC system over conventional ISAC systems. The main contributions of this paper are summarized as follows.
@@
  • In terms of sensing processing, we delve into the data-level and symbol-level fusion of the sensing data on high and low-frequency bands. Data-level fusion refers to the fusion of multiple target information estimated by sensing data, while symbol-level fusion refers to the fusion of received symbols with phase information [18, 19]. In the symbol-level fusion of sensing data, two challenges manifest: 1) The initial phase misalignment of the echo signals on high and low-frequency bands due to attenuation and radar cross section (RCS); 2) The fusion of the sensing data on high and low-frequency bands with different physical-layer parameters. To this end, this paper decomposes the sensing signal processing into signal preprocessing stage and sensing information fusion stage. The signal preprocessing stage fully leverage the sensing data from multiple antennas while aligning the initial phases of the echo signals on high and low-frequency bands using the cyclic cross-correlation (CCC) method, as detailed in Section III-A. In the sensing information fusion stage, this paper realizes the symbol-level fusion of the sensing data on high and low-frequency bands, as well as the range and velocity estimations of targets by rearranging the high and low-frequency feature vectors and using the premise of cyclic prefix (CP) dynamic adjustment, as detailed in Section III-B.
  • The theoretical boundaries of the proposed CA-enabled MIMO-OFDM ISAC signal are derived. Specifically, mutual information (MI) is adopted as the performance metric for communication, while the Cramér-Rao lower bound (CRLB) is utilized as a typical performance metric of sensing. Based on the proposed ISAC system, we derive the closed-form communication MI under the frequency-selective fading MIMO-OFDM channel model and sensing CRLBs for range and velocity estimations under the MIMO-OFDM sensing signal model.
@@ II System Model
We consider a CA-enabled MIMO-OFDM ISAC base station (BS) for DL communication and sensing potential targets as shown in Fig. 1. The ISAC BS is equipped with $N_{\text{T}}+N_{\text{R}}$ antennas, where $N_{\text{T}}$ transmit antennas in transmitter (Tx) simultaneously serve $U$ user equipments (UEs) with $N_{\text{U}}$ antennas and detect $I$ potential targets. $N_{\text{R}}$ receive antennas in receiver (Rx) are equipped to receive the echo signals reflected by targets [20]. The spacing of antennas in BS is denoted by $d_{\text{r}}$ [21]. The ISAC BS can transmit DL data information while simultaneously emitting sensing signals for target sensing. Data echo signals are used for UE sensing, and sensing echo signals for target sensing. Since the sensing processing methods are identical, this paper focuses on the processing of sensing echo signals.
@@ II-A CA-enabled MIMO-OFDM ISAC Signal
represents the data for all UEs on $N_{b}$ subcarriers during $M_{b}$ OFDM symbol times, where $\mathbf{S}_{n}^{b}=\left[\mathbf{s}_{n,1}^{b},\mathbf{s}_{n,2}^{b},\cdots,\mathbf{s}_{n,U}^{b}\right]\in\mathbb{C}^{M_{b}\times U}$ indicates the data for all UEs on the $n$-th subcarrier during $M_{b}$ symbol times with $\mathbf{s}_{n,u}^{b}\in\mathbb{C}^{M_{b}\times 1}$ being the specific symbol data vector for the $u\in\{1,2,\cdots,U\}$-th UE [22]. There is a trade-off between the randomness of the communication signal and the certainty of the sensing signal. To guarantee the balanced ISAC performance, this paper adopts the low complexity data independent precoding (DIP) scheme proposed in [23], which can be applied to various ISAC systems. Specifically, an ISAC precoding matrix [22, 23]
@@
$x^{b}\left(k,n^{\prime},m\right)=\frac{1}{N_{b}}\sum_{n=0}^{N_{b}-1}\mathbf{x}_{k}^{b}\left(n,m\right)e^{j2\pi\frac{n^{\prime}n}{N_{b}}},$ (4)
where $\mathbf{x}_{k}^{b}\left(n,m\right)$ denotes the $\left(n,m\right)$-th element of $\mathbf{x}_{k}^{b}=[\mathbf{X}^{b}]_{(k-1)N_{b}+1:kN_{b},:}$. $n^{\prime}\in\{0,1,\cdots,N_{b}-1\}$ represents the index of subcarriers in time-domain and $k\in\{0,1,\cdots,N_{\text{T}}-1\}$ represents the index of transmit antenna. A CP with size $N_{\text{s}}^{b}$ greater than the maximum delay caused by target reflection is added to avoid inter-symbol interference. After digital-to-analog conversion (DAC), the analog signals of the $1$-st and $2$-nd CCs are up-converted using different local oscillators (LOs), mixed, and transmitted via a linear high power amplifier (HPA), which avoids the nonlinear effects.
@@
$x(t)_{k}=\sum_{b=1}^{\mathcal{B}}\sum_{m=0}^{M_{b}-1}\sum_{n=0}^{N_{b}-1}\mathbf{x}_{k}^{b}\left(n,m\right)e^{j2\pi t(f_{\mathrm{C}}^{b}+n\Delta f^{b})}\text{rect}\left(\frac{t-mT^{b}}{T^{b}}\right),$ (5)
@@ II-B Communication Model
In the communication channel model, since the high and low-frequency bands are processed independently without fusion, the differences in attenuations caused by high and low-frequency bands are not considered.
@@ II-B1 Communication channel model
is the channel matrix of the $l\in\{0,1,\cdots,L-1\}$-th delay path for the $u$-th UE with $h_{k,z}^{u,l,b}\sim\mathcal{CN}(0,1)$ being the corresponding channel element with independent identically distributed (i.i.d.) Rayleigh fading coefficient [25, 22]. Therefore, the MIMO communication channel matrix in time-domain for the $u$-th UE on the $n$-th subcarrier during the $m$-th OFDM symbol time is expressed as [26, 24]
@@ II-B2 Received communication signal
At the DL preamble+pilots information period, pilots are used to obtain channel state information (CSI) estimation [27, 28]. At the DL data information period, the $u$-th UE receives the DL communication signal, and undergoes down-conversion, analog-to-digital conversion (ADC), CP removal, DFT, and other operations to get a received data $\mathbf{Y}_{u}^{\mathrm{C}}$, which is expressed as
@@
Since linear HPA is considered in this paper, it is proposed that the communication processing of each CC of the proposed ISAC system is similar to that of the traditional MIMO-OFDM communication system, and is independent for each UE [29, 24]. In terms of communication performance in CA-enabled MIMO-OFDM ISAC system, the closed-form communication MI is derived in Section IV-A.
@@ II-C Sensing Model
To better reflect practical scenarios, we consider the factors that directly or indirectly influence the high and low-frequency aggregation in sensing signal processing, including phase alignment errors caused by attenuation differences, physical parameter variations, and noise variance discrepancies [30]. To the best of our knowledge, this is the first work in CA-enabled ISAC research to consider phase alignment errors and noise variance discrepancies, making it more applicable to practical applications.
@@ II-C1 Sensing channel model
We assume that the radial velocity of the $i$-th potential target is $v_{i,0}$ and the range to BS is $r_{i,0}$. Meanwhile, the angle of arrive (AoA) and the angle of departure (AoD) between the $i$-th target and BS are identical and are denoted by $\theta_{i,\text{Rx}}=\theta_{i,\text{Tx}}$. The transmit and receive steering vectors are expressed as (9) and (10), respectively.
$\mathbf{a}_{\text{Rx}}(\theta_{i,\text{Rx}})$ $=\left[e^{j2\pi p(\frac{d_{\text{r}}}{\lambda^{b}})\sin(\theta_{i,\text{Rx}})}\right]^{\text{T}}|_{p=0,1,\cdots,N_{\text{R}}-1},$ (9)
$\mathbf{a}_{\text{Tx}}(\theta_{i,\text{Tx}})$ $=\left[e^{j2\pi k(\frac{d_{\text{r}}}{\lambda^{b}})\sin(\theta_{i,\text{Tx}})}\right]^{\text{T}}|_{k=0,1,\cdots,N_{\text{T}}-1},$ (10)
@@
The sensing channel model on the $n$-th subcarrier during the $m$-th OFDM symbol time in the $b$-th CC is expressed as [28, 31]
$\mathbf{H}_{m,n}^{b}=\sum_{i=1}^{I}\left[\begin{array}[]{l}\kappa_{i,S}^{b}e^{j2\pi\left(f_{i,\text{s}}^{b}mT^{b}-n\Delta f^{b}\tau_{i,0}\right)}\\ \mathbf{a}_{\text{Rx}}(\theta_{i,\text{Rx}})\mathbf{a}_{\text{Tx}}^{\text{T}}(\theta_{i,\text{Tx}})\end{array}\right],$ (11)
where $\mathbf{H}_{m,n}^{b}\in\mathbb{C}^{N_{\text{R}}\times N_{\text{T}}}$ and $\kappa_{i,S}^{b}=\sqrt{\frac{(\lambda^{b})^{2}}{(4\pi)^{3}r_{i,0}^{4}}}\beta_{i,S}^{b}$ is the attenuation between the $i$-th target and BS with $\beta_{i,S}^{b}\sim\mathcal{CN}(0,\sigma_{\beta_{i,S}^{b}}^{2})$ being the $i$-th target’s RCS [32], which characterizes the difference in the initial phase of high and low-frequency bands; $f_{i,\text{s}}^{b}=\frac{2f_{\mathrm{C}}^{b}v_{i,0}}{c}$ and $\tau_{i,0}=\frac{2r_{i,0}}{c}$ are the Doppler frequency shift and delay, respectively.